# Scalable and Domain-General Abstractive Proposition Segmentation

### v1

@@ 6 The propositions API
We showed that our student models resolve two issues with the commonly used few-shot prompting approach: under-extraction (low $RL_{r}$) and cost. As part of this paper, we release a proposition segmentation API based on the Gemma 7B student model trained from Gemini Pro data (grouped propositions version). [^7] Link will be provided soon. We invite researchers that require proposition segmentation on input text to try out our models instead of few-shot prompting LLMs.
@@ 7 Conclusion
We define the abstractive propositions segmentation task more formally by specifying the desired properties of propositions and present a suite of automatic evaluation metrics that allow us to measure different dimensions of quality. While previous work often uses few-shot prompting, we show that supervision from existing datasets yields significant quality improvement. We then propose a distillation approach for training scalable and domain-general models that get on-par results with the teachers (and sometimes even better). We release an API based on Gemma 7B student models and invite researchers to use that instead of few-shot prompting LLMs.
@@ 8 Limitations
Appendix A Few-shot Prompting Example for Synthetic Multi-Domain Text Generation
@@
Appendix B Hyper-parameters
@@
Appendix C Aligning Propositions with Sentences and Filtering Problematic Examples
As explained in §5.1, we align propositions in the ROSE dataset with their corresponding sentences, and filter problematic examples. In particular, we follow these steps:
@@ Appendix C Aligning Propositions with Sentences and Filtering Problematic Examples
Appendix D Full Results on All Datasets
In this section, we show all the results reported in §5 plus two additional sets of results. Table 8, 9, and 10 show the full results.
@@ Appendix D Full Results on All Datasets
Appendix E Few-shot Models Recall Issues

### v2

@@ 7 The propositions API
We showed that our student models resolve two issues with the commonly used few-shot prompting approach: under-extraction (low $RL_{r}$) and cost. As part of this paper, we release the Gemma-APS API on Hugging Face based on Gemma 1 2B G and Gemma 1 7B G student models trained from Gemini Pro data (grouped propositions version). We invite researchers that require proposition segmentation on input text to try out our models instead of few-shot prompting LLMs.
@@ 8 Conclusion
We define the abstractive proposition segmentation task more formally by specifying the desired properties of propositions and present a suite of automatic evaluation metrics that allow us to measure different dimensions of quality. While previous work often uses few-shot prompting, we show that supervision from existing datasets yields significant quality improvement. We then propose a distillation approach for training scalable and domain-general models that get on-par results with the teachers (and sometimes even better). We release an API based on Gemma 7B student models and invite researchers to use that instead of few-shot prompting LLMs.
@@ 9 Limitations
We used NLI as the backbone to our metrics. We note that NLI as a task is not fully solved, and there are some levels of disagreement in human annotation Pavlick and Kwiatkowski (2019); Weber-Genzel et al. (2024). However, we showed strong correlations between human judgements and the defined metrics. In addition, NLI computation is done using a fine-tuned language model, so it is not very lightweight. However, the metric computation usually needs to be done on a small dataset.
@@
Appendix A Metric Calculation Examples
We show examples of how metrics are calculated in Table 6. In each row, we show an example and the expected and calculated metrics for it. We also mention which propositions property is mainly measured by the metric. Finally, we provide an explanation about how the property affects the score.
Metric Property Example Expected Score Calculated Score Explanation
Reference-less precision Supported Passage = "The price of the books are all less than ten dollars, and they download before you can get up for a cup of coffee." – Predicted Propositions = ["The books download before you can get up for a cup of coffee."] 1 0.9999 The predicted proposition is entailed by the passage.
Reference-less recall Comprehensive Passage = "The price of the books are all less than ten dollars, and they download before you can get up for a cup of coffee." – Predicted Propositions = ["The price of the books are all less than ten dollars."] 0 0 The predicted propositions do not cover all the information in the passage.
Reference-based precision Self-contained Predicted Propositions = ["The price of the books are all less than ten dollars.", "They download before you can get up for a cup of coffee."] – Gold Propositions = ["The price of the books are all less than ten dollars.", "The books download before you can get up for a cup of coffee."] 0.5 0.5 The second predicted proposition is not as self-contained as the second gold proposition ("They" vs "The books"). Therefore, the second predicted proposition should get a score of 0 when calculating reference-based precision.
Reference-based recall Self-contained Predicted Propositions = ["The price of the books are all less than ten dollars.", "They download before you can get up for a cup of coffee."] – Gold Propositions = ["The price of the books are all less than ten dollars.", "The books download before you can get up for a cup of coffee."] 0.5 0.5 The second gold proposition is more self-contained than the second predicted proposition ("The books" vs "They"). Therefore, the second gold proposition should get a score of 0 when calculating reference-based recall.
Reference-based precision Atomic Predicted Propositions = ["The price of the books are all less than ten dollars. They download before you can get up for a cup of coffee."] – Gold Propositions = ["The price of the books are all less than ten dollars.", "The books download before you can get up for a cup of coffee."] 0 0 The predicted proposition is not as atomic as any of the gold proposition. Therefore, it gets a score of 0 when calculating reference-based precision.
Reference-based recall Atomic Predicted Propositions = ["The price of the books are all less than ten dollars. They download before you can get up for a cup of coffee."] – Gold Propositions = ["The price of the books are all less than ten dollars.", "The books download before you can get up for a cup of coffee."] 0 0 The gold propositions are more atomic than the predicted proposition. Therefore, they both get a score of 0 when calculating reference-based recall.
Table 6: Examples with expected and calculated metrics. For each example, we provide the propositions property that is mainly measured by the metric. In addition, we explain how the property affects the score.
Appendix B Few-shot Prompting Example for Synthetic Multi-Domain Text Generation
@@
Appendix C Hyper-parameters
@@
Appendix D Pre-processing the ROSE Dataset, Aligning Propositions with Sentences, and Filtering Problematic Examples
We pre-process the dataset to improve its quality based on the following steps:
A) In some cases, ACUs end with a space before the period. We remove the extra space. Additionally, some ACUs do not end with a period (and do not end with “…” either). In these cases, we add a period to the ACU.
B) For each sentence, the ROSE dataset annotators first write an ACU consisting the main information from the subject of the main clause. Then, they add one ACU for each additional information in the sentence by adding minimal necessary information to the original ACU Liu et al. (2023). In some cases, the original ACU is exactly repeated in other ACUs. In these cases, we removed the first ACU as they are often very short and not very informative. For example, the ACU “Many seals are shot” is removed because we also have another ACU “Many seals are shot to death for their fur”.
C) As explained in §5.1, we align propositions in the ROSE dataset with their corresponding sentences, and filter problematic examples. In particular, we follow these steps:
@@
Appendix E Full Results on All Datasets
In this section, we show all the results reported in §5 plus two additional sets of results. Table 10, 11, and 12 show the full results.
@@ Appendix E Full Results on All Datasets
Appendix F Few-shot Models Recall Issues